OpenStudy (anonymous):

lincoln and gemma are looking at the equation sqrt (2x + 3) = x. Lincoln says that the solution is extraneous. Gemma says the solution is non extraneous Whos correct? MEDAL WILL BE GIVEN

OpenStudy (anonymous):

@Becki

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

Square both sides and set it all equal to zero. Factor. Then solve. You'll get 2 answers. Put each one in for x in the original equation. Are either of your equations with each of your 2 answers true? If yes, then nonextraneous. If no, then extraneous. :)

OpenStudy (anonymous):

wait so 2x + 3 = 0 and x = 0?

OpenStudy (anonymous):

@Becki

OpenStudy (anonymous):

No. [sqrt (2x+3)]^2=x^2 So if you square the square root of something you get...?? :) Think of (sqrt2)*(sqrt2)

OpenStudy (anonymous):

(2x + 3)(2x+3)

OpenStudy (anonymous):

Not exactly. It would be sqrt(2x+3)sqrt(2x+3) which would just equal (2x+3). So on one side you have (2x+3) and on the other x^2. \[2x+3=x ^{2}\] Now put everything on the same side, setting it equal to 0. \[x ^{2}-2x-3=0\] and factor (x- ?)(x+ ?)=0

OpenStudy (anonymous):

ehh...im lostt

OpenStudy (anonymous):

why is it not +3

OpenStudy (anonymous):

Because we moved everything to the side with the x^2. 2x+3=x^2 2x =x^2-3 0 =x^2-2x-3 I just turned it around so the zero was on the right.

OpenStudy (anonymous):

2x+3=x^2 2x+3-3=x^2-3 (subtract 3 from both sides) 2x-2x=x^2-2x-3 (subtract 2x from both sides)

OpenStudy (anonymous):

do u think u can do it with the equation button below!

OpenStudy (anonymous):

this 2x + 3 = x^2 is confusing now. Where did u get that

OpenStudy (anonymous):

\[2x+3-3=x ^{2}-3\] \[2x-2x=x ^{2}-2x-3\] \[0=x ^{2}-2x-3x \] Turned around: \[x ^{2}-2x-3=0\]

OpenStudy (anonymous):

2x+3=x^2 came from squaring both sides. If you take \[\sqrt{2x+3}\times \sqrt{2x+3}\] you get \[2x+3\]

OpenStudy (anonymous):

ok got that part !

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

I say lincon is correct

OpenStudy (anonymous):

I just need the steps on how to cause they want me to type it

OpenStudy (anonymous):

ah.... then nvm haha

OpenStudy (anonymous):

id go with hero on this

OpenStudy (anonymous):

@Luigi0210 @galacticwavesXX @recon14193

OpenStudy (agent0smith):

\[\Large x ^{2}-2x-3=0\] do you understand up to here?

OpenStudy (anonymous):

Yes!

OpenStudy (agent0smith):

K... can you factor it? you want two numbers that multiply to make the -3, and add up to -2

OpenStudy (anonymous):

I cant remember how to factor to well. which two numebers do u use

OpenStudy (agent0smith):

what numbers multiply to make -3? list them (there are not many)

OpenStudy (anonymous):

1 and -3 .. and idk I only came up with one

OpenStudy (anonymous):

when you have \[x^2-2x-3=0\] this polynomial can be factored or simplified \[x^2-2x-3=0 \rightarrow (x-3)(x+1)=0\] -3 + 1 = -2 that is where you get the -2x in the polynomial and -3*1 = -3 that is where the -3 came from x*x =x^2 and putting it together gives you the polynomial you understand that so far?

OpenStudy (anonymous):

yess!

OpenStudy (anonymous):

squaring both sides is the standard way of solving this kind of problem

OpenStudy (anonymous):

what happened to the sqrt in your solving method

OpenStudy (anonymous):

haha it's all good

OpenStudy (anonymous):

@crystelle x=3 and x=-1 because (x-3)=0 & (x+1)=0, you solve for x

OpenStudy (anonymous):

now that you know what the solutions for x are plug in x into the original equation to check if the solutions are extraneous or not \[\sqrt{2x+3}=x\] plug in x = 3 and x =-1 at separate times \[\sqrt{2(3)+3}=3 \rightarrow \sqrt{6+3}=3 \rightarrow \sqrt{9}=3 \rightarrow 3=3\] \[\sqrt{2(-1)+3}=3 \rightarrow \sqrt{-2+3}=3 \rightarrow \sqrt{1}=3 \rightarrow 1\neq3\] Even thought the solution x = 3 satisfies the equation, x = -1 doesn't so the solutions are extraneous Therefore lincoln is correct

OpenStudy (anonymous):

You understand now how i solved and got the answer?

OpenStudy (anonymous):

wait it should be 1 doesn't equal -1 but Lincoln is still correct sorry for the error

OpenStudy (anonymous):

im reviewing it!

OpenStudy (anonymous):

Ok im see it now

OpenStudy (anonymous):

wait in the problem It says is gemme or lincoln correct or are they both correct